1. Using the Weiszacker’s formula, calculate the binding energy per nucleon for 235


92 U and


its two fission products: 149


60 Nd and 86


32Ge.



1
Expert's answer
2021-12-27T08:03:01-0500

"\omega=\frac EA=\frac{c^2}A(Zm_p+Nm_n-m),"

"\omega_{U}=7.59~\frac{MeV}A,"

"\omega_{Nd}=8.26~\frac{MeV}A,"

"\omega_{Ge}=8.36~\frac{MeV}A."